OpenStudy (anonymous):

What is x-2y=1 in slope intercept form?

OpenStudy (anonymous):

translate as "solve for \(y\)"

OpenStudy (anonymous):

It's -1/2

OpenStudy (anonymous):

takes two steps a) subtract \(x\) from both sides b) divide both sides by \(-2\)

OpenStudy (anonymous):

actually it is not \(-\frac{1}{2}\)

OpenStudy (anonymous):

I got y= -1/2 - x

OpenStudy (anonymous):

y=1/2x−1/2

OpenStudy (anonymous):

x=2y+1

OpenStudy (anonymous):

Hmm... okay. I'm not quite sure how you got there, but I'm really bad at this stuff :P

OpenStudy (anonymous):

https://www.mathway.com/

OpenStudy (anonymous):

it helps alot

OpenStudy (anonymous):